Two complementary angles are in ratio 1:5. Fine the two angles

Answer by math_tutor2020(3816)


One angle is x, and the other is 5x
They form the ratio x:5x which reduces to 1:5
x is some positive real number

Those two angles are complementary, meaning they add to 90 degrees
x+5x = 90
6x = 90
x = 90/6
x = 15

This leads to 5x = 5*15 = 75

The two angles are 15 degrees and 75 degrees.